What average net force is needed to accelerate a 1380 kg car to a speed of 25 m/s in a time of 7 s?

N
Answer #1

Solution:

To calculate the force we need acceleration,

Using v = u + a * t

=> 25 = 0 + a * 7

=> a = 25 / 7 m/s^2

F = m * a

=> F = 1380 * 25 / 7

=> F = 4928.57 N

please comment for Queries.

Please rate thanks.